OpenStudy (anonymous):

hi i need help... again: Identify the slope and y-intercept of the inequality 9x + 3y > 6. slope = 3, y-intercept = 2 slope = -3, y-intercept = 2 slope = -3, y-intercept = 6 slope = -1/3, y-intercept =2 and this one: Renee needs to burn at least 300 calories at the gym today. She can burn 25 calories per minute on the treadmill (t) and 10 calories per minute on the elliptical trainer (e). Which inequality below can be used to find the possible combinations of time she could s

OpenStudy (anonymous):

oh did it off? alright, hold on.

OpenStudy (anonymous):

cut off* curse this lag.

OpenStudy (anonymous):

she could spend on each machine and still meet her goal of burning at least 300 calories? 25t + 10e >= 300 25t - 10e >= 300 25t + 10e <= 300 25t - 10e <= 300 anyways, if you could please tell me the correct answer and how you got it, it would be greatly appreciated

OpenStudy (anonymous):

45>85

OpenStudy (anonymous):

For the second question: She wants to burn AT LEAST 300 calories! burns 25 calories/min on treadmill: 25t burns 10 calories/min on the elliptical trainer: 10e so 25t and 10e have to be equal to AT LEAST 300! Hope I helped!

OpenStudy (anonymous):

For the first equation you need to change 9x + 3y > 6 to slope-intercept form: 9x + 3y > 6 -9x 3y > -9x + 6 /3 y > -3x + 3 is the final equation. Remember that y = mx + b where m=slope and b = y-intercept!

OpenStudy (anonymous):

I hope I helped you understand!!

OpenStudy (anonymous):

yeah, thanks jazy

OpenStudy (anonymous):

closing it now.